OpenStudy (khally92):

differenciate using 1st principle square root of 2x f(x+h)-f(x)/h

OpenStudy (khally92):

i got 1/x^1/2 using power rule but not the same as when using 1st principle

OpenStudy (khally92):

is it that difficult to solve..

OpenStudy (anonymous):

\[\lim_{h \to 0}\frac{f(x+h)-f(x)}{h}=\lim_{h \to 0}\frac{\sqrt{2(x+h)}-\sqrt{2x}}{h}\] is a start then rationalize the numerator, cancel one factor of h and replace it by zero to take the limit

OpenStudy (khally92):

not working i got 1/squareroot of 2x..u gotta help pls

OpenStudy (khally92):

after rationalizing i got 2x-2h+2x/denominator which after taking limit becomes squareroot of 2x. and the numeratot 2x cancels -2x leaving

OpenStudy (anonymous):

\[\frac{f(x+h)-f(x)}{h}=\frac{\sqrt{2(x+h)}-\sqrt{2x}}{h}\] \[=\frac{\sqrt{2x+2h}-\sqrt{2x}}{h}\] \[=\frac{\sqrt{2x+2h}-\sqrt{2x}}{h}\times\frac{\sqrt{2x+2h}+\sqrt{2x}}{\sqrt{2x+2h}+\sqrt{2x}}\] \[=\frac{2x+2h-2x}{h(\sqrt{2x+2h}+\sqrt{2x})}\] \[=\frac{2h}{h(\sqrt{2x+2h}+\sqrt{2x})}\] \[=\frac{2}{(\sqrt{2x+2h}+\sqrt{2x})}\] now take the limit by replacing h by zero

OpenStudy (khally92):

thats 2/2x+squareroot of 2.. am i 2 solve futher

OpenStudy (khally92):

thats 1/xsquareroot of 2.

OpenStudy (khally92):

buh its not thesame as when using power rule which gives ue x^-1/2

OpenStudy (khally92):

thnks ill jus leave the answer like that thnks alot man..